# Imitation from Observations with Trajectory-Level Generative Embeddings

### v1

@@ 5.1 Experimental Setup
Environments and Datasets
We evaluate our method on offline learning-from-observations (LfO) benchmarks built on D4RL, covering both locomotion and manipulation domains. The locomotion tasks are based on MuJoCo, including hopper, halfcheetah, walker2d, and ant, while the manipulation tasks are based on Adroit, including pen, door, and hammer. For MuJoCo, the offline suboptimal dataset $\mathcal{D}_{\mu}$ contains approximately 1M transitions drawn from the D4RL random or medium datasets, following established LfO protocols used in SMODICE (sikchi2024dual) and DILO (yan2024offline). For Adroit, $\mathcal{D}_{\mu}$ consists of human and cloned trajectories, again adhering to the standard benchmark protocol.
Following prior work, we consider two mixture settings that differ in the proportion of expert transitions included in $\mathcal{D}_{\mu}$: (i) expert, in which $\mathcal{D}_{\mu}$ contains 200 expert trajectories, and (ii) few-expert, in which $\mathcal{D}_{\mu}$ contains 30 expert trajectories. For the detailed mixture strategy, we refer to Appendix B.
Expert data
All methods are additionally provided with an observation-only demonstration $\mathcal{D}_{E}$ consisting of a single expert trajectory, which supplies the expert state sequence required by learning-from-observations (LfO) algorithms.
Baselines and Evaluation
We compare against recent offline learning-from-observations (LfO) methods, including SMODICE ma2022versatile, PW-DICE yan2024offline, and DILO sikchi2024dual, which have been shown to outperform earlier approaches such as LobsDICE kim2022lobsdice and DemoDICE kim2021demodice. We additionally include trajectory-level behavior cloning (BC) as a baseline, trained on the suboptimal data using a diffuser (janner2022planning) without reward guidance. Each configuration is evaluated with three random seeds, and we report the mean and standard deviation of normalized D4RL scores across seeds for all tasks and methods. We provide implementation details and hyperparameter settings in Appendix A, along with detailed training dynamics visualizations in Appendix D.
Table 1: Main Results. TGE (combined with IQL or ReBRAC, highlighted in cyan) is compared against state-of-the-art offline LfO baselines. The results demonstrate that TGE achieves superior performance. We report the mean and standard deviation for each result with 3 random seeds. Results that show statistically significant improvements over the second-best method (t-test) are highlighted in bold.
@@ 5.2 Experimental Results
As shown in Table 1, TGE combined with existing offline RL backbones consistently matches or outperforms all baselines. The performance gains are most pronounced in the medium+few-expert and medium regimes across all environments (hopper, halfcheetah, walker2d, and ant).
@@
In these settings, where the offline dataset contains limited expert-level data, occupancy-matching methods such as DILO (sikchi2024dual) and PW-DICE (yan2024offline) suffer substantial performance degradation and often fail to recover a functional policy. In contrast, TGE-based reward estimation, when combined with standard offline RL backbones, maintains robust performance across these regimes, resulting in a clear margin over prior approaches. These empirical findings support our analysis in Section 4.4, which suggests that occupancy-matching objectives become less informative under severe support mismatch between $\mathcal{D}_{\mu}$ and $\mathcal{D}_{E}$. By constructing rewards based on geometric proximity in the latent space of the generative planner, rather than direct support overlap, TGE is able to recover meaningful reward signals even when the support mismatch is large, thereby bridging the gap where occupancy ratio–based methods tend to fail.
Performance on Adroit Manipulation Benchmarks
@@
In the high-dimensional Adroit domain of dexterous hand manipulation, we observe a trade-off between handling human-collected and generated data. When combined with downstream offline RL backbones, TGE-based reward estimation performs particularly well on the human+expert datasets, establishing a new state of the art by consistently outperforming DILO (sikchi2024dual) and PW-DICE (yan2024offline) across all tasks. These results highlight the robustness of TGE to diverse and potentially non-Markovian noise present in human demonstrations.
@@
On the cloned+expert datasets, performance is more mixed. While TGE achieves the highest score on pen, it underperforms DILO on door and hammer. A possible explanation is that cloned data is generated by a relatively consistent policy, under which density-ratio–based methods such as DILO may be better able to exploit the resulting structural overlap between expert and generated data distributions.
@@
Nevertheless, the strong performance of TGE-based method on human data underscores its practical relevance for real-world settings, where data collection is often noisy and less structured.
(a) Reward Distribution Comparison
(b) Performance
Figure 3: Ablation of the Reward Signal Density. (Top) Walker2d-Medium. (Bottom) Hopper-Random. Distributions of synthetic rewards assigned to expert (red) and suboptimal (blue) data. Compared to Gaussian-kernel-based reward estimation, the Laplacian kernel preserves a heavier-tailed reward signal and is associated with improved performance and training stability.
5.3 Ablation Studies
Choice of Kernels for Reward Estimation
We investigate the choice of kernel function by analyzing the resulting reward distributions and their associated training dynamics. Figure 3(a) visualizes the density of synthetic rewards assigned to expert and suboptimal trajectories in the offline dataset under Laplacian and Gaussian kernels, respectively.
As shown in Figure 3(a, right), the Gaussian kernel produces a reward distribution that is highly concentrated in narrow regions: expert rewards collapse near $1.0$, forming a sharp peak, while suboptimal trajectories cluster nearby. As a result, meaningful behavioral differences correspond to only minor reward variations, which limits the clarity of the learning signal. In contrast, the Laplacian kernel (Figure 3(a, left)) yields a broader, heavy-tailed reward distribution. Its linear log-decay better preserves the relative ordering of trajectories in the embedding space, improving reward separability. This enhanced distinguishability facilitates more accurate value estimation by the critic and leads to more stable and optimal learning curves, as shown in Figure 3(b).
Impact of Horizons in the Generative Planner
We further examine how the length of temporal context used by the generative planner influences reward estimation quality and, in turn, the performance of the downstream offline RL policy. Specifically, we vary the horizon length $H$ employed during planner training and embedding computation. As shown in Figure 4, using a single-step encoding (H=1H{=}1) leads to a substantial degradation in the final policy performance, highlighting the importance of incorporating temporal context along the trajectory. Increasing $H$ yields nearly consistent performance improvements across tasks, with gains saturating around H=32H{=}32. This trend suggests that short-to-mid temporal windows are sufficient to capture essential dynamic information, while longer horizons provide diminishing returns and may introduce unnecessary computation or over-smoothing effects. Based on these observations, we adopt H=32H{=}32 as the default setting, which provides a strong and stable trade-off across environments.
We provide additional ablation studies on the kernel temperature and the number of nearest neighbors in Appendix C.
@@ 6 Conclusion
In this work, we introduced Trajectory-level Generative Embeddings (TGE), a novel framework for offline imitation learning from observations that addresses the structural limitations of distribution-matching approaches. By shifting the paradigm from density ratio estimation to generative metric learning, TGE overcomes the support coverage assumption that hampers prior methods. We demonstrated that encoding trajectories into the latent manifold of a temporal diffusion model allows for the construction of a dense, informative reward signal even when the offline data is distributionally disjoint from the expert.
These findings suggest that generative models can serve as powerful tools for structuring the state space in reinforcement learning, going beyond their traditional roles in planning and synthesis. Future work could explore extending this manifold-based reward framework to online fine-tuning settings or leveraging larger-scale video diffusion models to enable imitation from cross-embodiment observations.
@@ A.1 Hyperparameters
Kernel type Laplacian
@@
Dropout rate 0 (Mujoco) / 0.1(Adriot)
Table 4: Hyperparameter Settings for ReBRAC. For the actor BC coefficient, values $\{1.0,0.5,0.1,10.0\}$ are used for {Others, Hopper-random, HalfCheetah-random, Adroit-Cloned}, respectively. For the critic BC coefficient, values $\{1.0,0.5,0.0\}$ correspond to {Others, HalfCheetah-random, Adroit-Cloned}, respectively.
@@
Actor Learning rate 3e43e-4
Critic Learning rate 3e43e-4
@@
Layer normalization True (Mujoco) / False (Adriot)
@@
Actor BC coefficient ($\beta_{\text{actor}}$) 1.0, 0.5, 0.1, 10.0
Critic BC coefficient ($\beta_{\text{critic}}$) 1.0, 0.5, 0.0
@@ A.2 Implementation Details
We implement the TGE reward as the log-likelihood of an $m$-nearest-neighbor kernel density estimator (KDE) defined in the trajectory-level latent embedding space. All embeddings are first normalized to the unit hypersphere ($z\leftarrow z/\lVert z\rVert_{2}$). Distances are computed using Euclidean distance on the sphere, which is monotonic with cosine distance ($\lVert z_{i}-z_{j}\rVert_{2}=\sqrt{2-2\cos\theta}$).
For each offline transition, we retrieve its $m$ nearest expert neighbors in the embedding space (with m=10m{=}10 by default) and compute kernel values using a Laplacian kernel, $f=\exp(-d/\tau)$. We adopt this linear log-decay rather than a Gaussian kernel with quadratic decay, as the latter tends to produce overly sparse reward signals in high-dimensional embedding spaces and weak supervision for suboptimal data. The neighbor kernel values are then aggregated using a Log-Sum-Exp operation to produce the final reward, which can be interpreted as a robust soft minimum distance to the expert support.
Finally, we train offline policies using IQL and ReBRAC as drop-in backbones, relying solely on $\mathcal{D}_{\mu}$ and the learned TGE reward, without any online interaction.
@@ Appendix B Details of the Mixture Strategy for the Suboptimal Dataset
We provide a detailed breakdown of the data mixture strategy used to construct the suboptimal dataset $\mathcal{D}_{\mu}$. Following established protocols in offline imitation learning from observations introduced in [ma2022versatile, sikchi2024dual], we construct $\mathcal{D}_{\mu}$ by combining a large corpus of suboptimal transitions with a limited number of expert trajectories. Table 5 summarizes the detailed composition for all suboptimal datasets.
@@ Appendix C Additional Ablation Studies
C.1 Kernel Temperature
We investigate the sensitivity of our framework to the kernel temperature $\tau$, which regulates the decay rate of the Laplacian kernel and determines the sharpness of the reward signal in the latent space. Table 6 reports the normalized scores on Walker2d-medium and Hopper-random across a wide range of values $\tau\in\{0.1,0.5,1.0,2.0,5.0\}$.
As shown in the results, TGE demonstrates remarkable robustness to this hyperparameter. Despite varying $\tau$ by an order of magnitude, the performance fluctuations remain minimal. These findings confirm that the proposed trajectory-level embedding provides a well-structured metric space where the relative proximity of trajectories is preserved. Consequently, precise task-specific tuning of $\tau$ is unnecessary, and we adopt $\tau=1.0$ as the default to balance signal discriminativeness and smoothness.
Table 6: Ablation on Kernel Temperature ($\tau$). We compare normalized scores across varying temperatures. The method shows robust performance for $\tau\in[0.1,5.0]$.
Environment Kernel Temperature ($\tau$)
$\tau=0.1$ $\tau=0.5$ $\tau=1.0$ (Default) $\tau=2.0$ $\tau=5.0$
@@ C.1 Kernel Temperature
C.2 Number of Nearest Neighbors
@@
Table 7: Ablation on Number of Nearest Neighbors ($m$). We evaluate the impact of hyperparameter $m$ on the performance over Walker2d medium and Hopper random settings.
@@ Appendix D Training Dynamics Visualizations
(b) Walker2d-Medium
(c) HalfCheetah-Random
(d) Ant-Random
@@
(b) Walker2d-Random
(c) HalfCheetah-Random
@@
(f) Walker2d-Medium
(g) HalfCheetah-Medium

### v2

@@ 5.1 Experimental Setup
Environments and Datasets. We evaluate our method on offline learning-from-observations (LfO) benchmarks built on D4RL, covering both locomotion and manipulation domains. The locomotion tasks are based on MuJoCo, including hopper, halfcheetah, walker2d, and ant, while the manipulation tasks are based on Adroit, including pen, door, and hammer. For MuJoCo, the offline suboptimal dataset $\mathcal{D}_{\mu}$ contains approximately 1M transitions drawn from the D4RL random or medium datasets, following established LfO protocols used in SMODICE (Ma et al., 2022) and DILO (Sikchi et al., 2025). For Adroit, $\mathcal{D}_{\mu}$ consists of human and cloned trajectories, again adhering to the standard benchmark protocol.
Following prior work, we consider two mixture settings that differ in the proportion of expert transitions included in $\mathcal{D}_{\mu}$: (i) expert, in which $\mathcal{D}_{\mu}$ contains 200 unlabeled expert trajectories, and (ii) few-expert, in which $\mathcal{D}_{\mu}$ contains 30 unlabeled expert trajectories; and (iii) medium, in which no expert trajectories are added to $\mathcal{D}_{\mu}$. For the detailed mixture strategy, we refer to Appendix B.
Expert data. All methods are additionally provided with an observation-only demonstration $\mathcal{D}_{E}$ consisting of one single expert episode, which supplies the expert state sequences required by LfO algorithms.
Baselines and Evaluation. We compare against recent offline learning-from-observations (LfO) methods, including SMODICE Ma et al. (2022), PW-DICE Yan et al. (2024), and DILO Sikchi et al. (2025), which have been shown to outperform earlier approaches such as LobsDICE Kim et al. (2022) and DemoDICE Kim et al. (2021). We additionally include trajectory-level behavior cloning (BC) as a baseline, trained on the suboptimal data using diffuser (Janner et al., 2022) without reward guidance. We report the mean and standard deviation of normalized D4RL scores across five seeds for each tasks and methods. The pseudo code and hyperparameter are presented in Appendix A and the detailed training dynamics is visualized in Appendix D.
Table 1: Main Results. TGE (combined with IQL or ReBRAC, highlighted in cyan) is compared against state-of-the-art offline LfO baselines. The results demonstrate that TGE achieves superior performance. We report the mean and standard deviation for each result with five random seeds. The top-performing results and those with overlapping standard deviations are highlighted in bold.
@@ 5.2 Experimental Results
Performance on MuJoCo Locomotion Benchmarks. As shown in Table 1, TGE combined with existing offline RL backbones consistently matches or outperforms all baselines across MuJoCo locomotion environments.
In random+expert and medium+expert, where the offline buffer contains relatively sufficient expert-like coverage, TGE remains highly competitive and achieves strong performance. More importantly, the performance gains are most pronounced in random+few-expert and medium+few-expert, where expert trajectories are scarce and the offline dataset is dominated by suboptimal behaviors, inducing a substantially more challenging distributional mismatch with limited support overlap between $\mathcal{D}_{\mu}$ and $\mathcal{D}_{E}$. In these settings, occupancy-matching methods such as SMODICE (Ma et al., 2022), DILO (Sikchi et al., 2025) and PW-DICE (Yan et al., 2024) suffer substantial performance degradation and often fail to recover a functional policy. We observe similar robustness in the medium regime, where no unlabeled expert trajectories are mixed into $\mathcal{D}_{\mu}$, suggesting that TGE can still identify and exploit expert-like structure through trajectory-level generative geometry even under severe support mismatch.
Notably, in several settings TGE even surpasses the performance of the provided expert demonstrations. We hypothesize that this is because the single observation-only expert trajectory used for imitation may not represent the best expert behavior in the dataset. By measuring trajectory-level proximity in the diffusion embedding space, TGE can assign higher rewards to all expert-like trajectories and effectively recover high-quality behaviors present in the offline buffer.
@@
These empirical findings support our analysis in Section 4.5, which suggests that occupancy-matching objectives become less informative under severe support mismatch between $\mathcal{D}_{\mu}$ and $\mathcal{D}_{E}$. By constructing rewards based on geometric proximity in the latent space of the generative planner, rather than direct support overlap, TGE is able to recover meaningful reward signals even when the support mismatch is large, thereby bridging the gap where occupancy ratio–based methods tend to fail.
@@
Performance on Adroit Manipulation Benchmarks. In the high-dimensional Adroit domain of dexterous hand manipulation, we observe a trade-off between handling human-collected and generated data. When combined with downstream offline RL backbones, TGE-based reward estimation performs particularly well on the human+expert datasets, establishing a new state-of-the-art by consistently outperforming DILO (Sikchi et al., 2025) and PW-DICE (Yan et al., 2024) across all tasks. These results highlight the robustness of TGE to diverse and potentially non-Markovian noise present in human demonstrations.
@@
On the cloned+expert datasets, while TGE achieves the highest score on pen, it underperforms DILO on door and hammer. A possible explanation is that cloned data is generated by a relatively consistent policy, under which density-ratio-based methods such as DILO may be better able to exploit the resulting structural overlap between expert and generated data distributions.
@@
Nevertheless, the strong performance of TGE on human data underscores its practical relevance for real-world settings, where data collection is often noisy and less structured.
5.3 Impact of Horizons in the Generative Planner
@@
We examine how the length of temporal context used by the generative planner influences reward estimation quality and the performance of the downstream offline RL policy. Specifically, we vary the horizon length $H$ employed during planner training and embedding computation. As shown in Figure 3, using a single-step encoding (H=1H{=}1) leads to a substantial degradation in the final policy performance, highlighting the importance of incorporating temporal context along the trajectory. Increasing $H$ yields nearly consistent performance improvements across tasks when $H$ is small, and saturating after certain horizon lengths. This trend suggests that short-to-mid temporal windows are sufficient to capture essential dynamic information. While longer horizons maintain consistent performance without significant degradation, they may introduce unnecessary computation or over-smoothing effects. Based on these observations, we adopt a conservative default of H=32H{=}32 for a unified and stable hyperparameter choice across all environments and configurations. Additional ablation studies on other aspects of our proposed method is deferred to Appendix C.
@@ 6 Conclusion
In this work, we introduced Trajectory-level Generative Embeddings (TGE), a novel framework for offline imitation learning from observations that addresses the structural limitations of distribution-matching approaches. By shifting the paradigm from density ratio estimation to generative representation learning, TGE overcomes the support coverage assumption that hampers prior methods. Motivated by the theoretically justified particle-based entropy estimation, we demonstrated that encoding trajectories into the latent space of a temporal diffusion model allows for the construction of a dense, informative reward signal even when the support mismatch between offline data and the expert is severe.
These findings suggest that generative models can serve as powerful tools for structuring the state space in RL, going beyond their traditional roles in RL-based planning and synthesis. Future work could explore extending this trajectory-level latent reward framework to online fine-tuning settings or leveraging larger-scale video diffusion models to enable imitation from cross-embodiment observations.
@@ A.1 Hyperparameters
Kernel type Logarithmic
@@
Dropout rate 0 (Mujoco) / 0.1(Adroit)
Table 4: Hyperparameter Settings for ReBRAC. For the actor BC coefficient, values $\{0.1,0.4,1.0\}$ are used for {Adroit-Cloned, Mujoco, Adroit-Human}, respectively.
@@
Actor Learning rate 5e45e-4
Critic Learning rate 5e45e-4
@@
Layer normalization True
Actor Dropout 0 (Mujoco) / 0.1(Adroit)
@@
Actor BC coefficient ($\beta_{\text{actor}}$) 0.1, 0.4, 1.0
Critic BC coefficient ($\beta_{\text{critic}}$) 1.0
@@ A.2 Implementation Details
We implement the TGE reward as the logarithmic distance to the nearest expert embeddings, consistent with the entropy minimization objective derived in Section 4.5. All embeddings are first normalized to the unit hypersphere ($z\leftarrow z/\lVert z\rVert_{2}$). Distances are computed using Euclidean distance on the sphere, which is monotonic with cosine distance ($\lVert z_{i}-z_{j}\rVert_{2}=\sqrt{2-2\cos\theta}$).
For each offline transition, we retrieve its $m$ nearest expert neighbors in the embedding space (with m=10m{=}10 by default) and compute the reward using a logarithmic kernel: $r=-\sum_{k=1}^{m}\log(1+d_{k}/\sigma)$, where $d_{k}$ is the distance to the $k$-th neighbor and $\sigma$ is a temperature scaling parameter. This reward signal is then normalized and used to label transitions in $\mathcal{D}_{\mu}$.
Finally, we train offline policies using IQL and ReBRAC as drop-in backbones, relying solely on $\mathcal{D}_{\mu}$ and the learned TGE reward, without any online interaction. Additionally, to ensure stable learning and balanced sampling across the diverse trajectories in the suboptimal dataset, we utilize the normalized TGE reward as a weighting factor for the behavioral cloning term in our ReBRAC backbone. We treat rewards as fixed confidence weights that scale the MSE penalty between the actor’s actions and the dataset actions. This ensures that the policy regularizes more strongly toward expert-labeled regions (high reward) while maintaining flexibility in suboptimal regions, effectively preventing the policy from collapsing toward poor-quality data points.
A.3 Computational Overhead
As noted in the experimental evaluation, TGE involves a multi-stage pre-processing pipeline before policy learning begins. We report the typical wall-clock times for these stages below:
Generative Training: Training the trajectory diffusion model on the suboptimal dataset $\mathcal{D}_{\mu}$ requires approximately 4-6 hours on a single L40S GPU. While this is higher than the zero-cost initialization of occupancy-matching baselines, it is a one-time offline cost that enables the extraction of temporally coherent features.
Reward Annotation: Computing embeddings and performing the top-$m$ nearest neighbor search for 1M transitions takes roughly 1 hour.
Total Training Efficiency: Although the pre-processing is more intensive than baselines, it enables the use of standard offline RL backbones without any generative sampling overhead during the RL loop. This makes the downstream policy training phase as efficient as standard RL on labeled data.
@@ Appendix B Details of the Mixture Strategy for the Suboptimal Dataset
We provide a detailed breakdown of the data mixture strategy used to construct the suboptimal dataset $\mathcal{D}_{\mu}$. Following established protocols in offline imitation learning from observations introduced in [Ma et al., 2022, Sikchi et al., 2025], we construct $\mathcal{D}_{\mu}$ by combining a large corpus of suboptimal transitions with a limited number of expert trajectories. Table 5 summarizes the detailed composition for all suboptimal datasets.
@@ Appendix C Additional Ablation Studies
C.1 Choice of Kernels for Reward Estimation
We investigate the choice of kernel function $f(d)$ by analyzing the resulting reward distributions and their associated training dynamics. Figure 4(a) visualizes the density of synthetic rewards assigned to expert and suboptimal trajectories in the offline dataset under Logarithmic and Gaussian kernels, respectively.
As shown in Figure 4(a), the Gaussian kernel produces a reward distribution that is highly concentrated in narrow regions: expert rewards collapse near $1.0$, forming a sharp peak, while suboptimal trajectories cluster nearby. As a result, meaningful behavioral differences correspond to only minor reward variations, which limits the clarity of the learning signal. In contrast, the Logarithmic kernel (Figure 4(a, left)) yields a broader, heavy-tailed reward distribution. The slow decay better preserves the relative ordering of trajectories in the embedding space, improving reward separability. This enhanced distinguishability facilitates more accurate value estimation by the critic and leads to more stable and optimal learning curves, as shown in Figure 4(b).
(a) Reward Distribution Comparison
(b) Performance
Figure 4: Ablation of the Reward Signal Density. (Top) Walker2d-Medium. (Bottom) Hopper-Random. Distributions of synthetic rewards assigned to expert (red) and suboptimal (blue) data. Compared to Gaussian-kernel-based reward estimation, the Logarithmic kernel preserves a heavier-tailed reward signal and is associated with improved performance and training stability.
C.2 Kernel Temperature
We investigate the sensitivity of our framework to the kernel temperature $\sigma$, which regulates the decay rate of the kernel and determines the sharpness of the reward signal in the latent space. Table 6 reports the normalized scores on Walker2d-medium and Hopper-random across a wide range of values $\sigma\in\{0.1,0.5,1.0,2.0,5.0\}$.
As shown in the results, TGE demonstrates remarkable robustness to this hyperparameter. Despite varying $\sigma$ by an order of magnitude, the performance fluctuations remain minimal. These findings confirm that the proposed trajectory-level embedding provides a well-structured metric space where the relative proximity of trajectories is preserved. Consequently, precise task-specific tuning of $\sigma$ is unnecessary, and we adopt $\sigma=1.0$ as the default to balance signal discriminativeness and smoothness.
Table 6: Ablation on Kernel Temperature ($\sigma$). We compare normalized scores across varying temperatures. The method shows robust performance for $\sigma\in[0.1,5.0]$.
Environment Kernel Temperature ($\sigma$)
$\sigma=0.1$ $\sigma=0.5$ $\sigma=1.0$ (Default) $\sigma=2.0$ $\sigma=5.0$
@@ C.2 Kernel Temperature
C.3 Number of Nearest Neighbors
@@
Table 7: Ablation on Number of Nearest Neighbors ($m$). We evaluate the impact of hyperparameter $m$ on the performance over Walker2d-medium and Halfcheetah-random settings.
@@ Appendix D Training Dynamics Visualizations
(b) HalfCheetah-Medium
(c) Walker2d-Medium
(d) Ant-Medium
(e) Hopper-Random
(f) HalfCheetah-Random
(g) Walker2d-Random
(h) Ant-Random
@@
(b) HalfCheetah-Random
(c) Walker2d-Random
@@
(f) HalfCheetah-Medium
(g) Walker2d-Medium